OpenStudy (anonymous):

If f(x) = -7x + 22 f(2)= f(x+1)=

OpenStudy (anonymous):

?

OpenStudy (anonymous):

f(x)= -7(2)+22 f(x)= -14+22 ----> f(x)= 8 f(x)= -7(x+1)+22 f(x)= -7x -7+22 ----> f(x)= -7x + 15 f(x) is simply y so when you have a number where the x would be you substitute it for X in the equation.

OpenStudy (anonymous):

^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^ What he said. Although, this is the history section. This question would go in the Math section. Click on "Find More Subjects" to get there. Hope this helps! @Toonez